OpenStudy (anonymous):

how do you know the asymptote of (1/4)^x

jimthompson5910 (jim_thompson5910):

as x gets larger and larger, (1/4)^x gets smaller and smaller and gets closer to 0

jimthompson5910 (jim_thompson5910):

so the horizontal asymptote of f(x) = (1/4)^x is y = 0

OpenStudy (anonymous):

@jim_thompson5910 so the asymptote is 0?

jimthompson5910 (jim_thompson5910):

y = 0 yes that's the horizontal asymptote

OpenStudy (anonymous):

but my teacher drew it, and the asymptote went past zero

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = 0 is the horizontal line through y = 0 ie it's the x axis

OpenStudy (anonymous):

Oh wait, you're correct. Thank you!

jimthompson5910 (jim_thompson5910):

so basically f(x) = (1/4)^x doesn't cross the x axis

jimthompson5910 (jim_thompson5910):

yw